OpenStudy (anonymous):

Hi I am 55 years old and I have had to back school I am trying to figure out fractions is there anyone out there that can help my test is Monday HELP

OpenStudy (anonymous):

If you have a concrete problem you need help with this site is real good for that, but for a loose subject like fractions I really recommend this site, it's excellent. http://www.purplemath.com/modules/

OpenStudy (anonymous):

thank you but will this get me to the answers i like right now?

OpenStudy (anonymous):

Like I said, if you got a specific problem/question, bring it on. If not, spend 20min on purplemath (fractions part) and you will have learned alot.

OpenStudy (anonymous):

I am having more problems then I can tell you, right now all I am trying to is get the answer to 81/4+121/2+91/3= I have not been in school since 1983

OpenStudy (anonymous):

sorry all i am trying to is figure out how to brake this down to the lowest terms

OpenStudy (anonymous):

Is it (81/4 + 121/2) all divided by 3?

OpenStudy (anonymous):

Your thought process should be this: How can I change these fractions so I can just add them together? The answer is this: The lowest common integer that binds 4 3 and 2. 4*3=12 3*4=12 2*6=12 So to keep the same value on these integers you must keep the ratio. Thereby you multiply the numbers like this: (81/4)*(3/3) (121/2)*(6/6) (91/3)*(4/4) See what I did?